# Deconfined quantum criticality of frustrated hard-core dipolar bosons

### v1

@@ Abstract
In this letter, we instead explore a frustrated dipolar Bose-Hubbard (FDBH) model in the hard-core limit in one dimension, where nearest-neighbor interactions arise from dipolar interactions [36]. We show that the FDBH model hosts DQCPs, aligning with recent theoretical proposals [13, 37, 38, 39, 40] that describe a continuous quantum phase transition from an antiferromagnetic (AFM) state to a valence bond solid (VBS) state. We employ the density matrix renormalization group (DMRG) technique [41, 42, 43, 44, 45] to investigate quantum phase transitions of FDBH model by varying nearest-neighbor interactions. Our study reveals the presence of three distinct phases, delineated by two BKT transitions and one DQCP through the fidelity susceptibility. Intriguingly, our proposal reveals that DQCPs emerge from the fusion of two BKT transitions, providing an essential insight into the understanding of both BKT transitions and DQCPs.
@@
by utilizing the transformations $\sigma^{x}_{r}=b^{\dagger}_{r}+b_{r},~{}~{}\sigma^{y}_{r}=i(b^{\dagger}_{r}-b_{r}),~{}~{}\sigma^{z}_{r}=1-2n_{r}$, with $(\sigma^{x}_{r},\sigma^{y}_{r},\sigma^{z}_{r})$ being Pauli matrices. Here, the particle occupation numbers are represented by spins. Specifically, $\sigma_{z}=-\frac{1}{2}$ denotes the down spin when $n_{j}=1$, and $\sigma_{z}=\frac{1}{2}$ signifies the up spin when $n_{j}=0$. Utilizing spin operators to explore the physics of the hard-core bosonic model is a straightforward and effective approach.
@@
When $V=0$, the system shown in Eq.(3) become the XY model incorporating both the nearest-neighbor and next-nearest-neighbor hopping terms. In particular, when $t_{2}=0$, the system’s ground state is characterized by the gapless XY (superfluidity (SF) in bosonic language) phase (see Appendix A for details). Upon increasing $t_{2}$, the system undergoes a phase transition from the XY phase to the VBS (bond-ordered-wave (BOW) in bosonic language) phase [46, 47, 48, 49]. To verify if the transition between the XY phase and the VBS phase persists in the presence of interaction $V$, we calculate the ground-state phase diagram of the FDBH model described in Eq.(2) varying parameters $t_{2}/t_{1}$ and $V/t_{1}$ in terms of DMRG method. During DMRG simulations, we keep 500 matrix states and enforce $n_{\text{max}}=1$ per site. The full phase diagram is shown in Fig.2, which is obtained from the fidelity susceptibility [50, 51, 52, 53, 54, 55, 56, 57],
@@
where, $\chi_{\infty}$ represents the fidelity susceptibility in the thermodynamic limit, $\lambda_{c}$ is the critical point, and $a$ is a positive nonuniversal constant serving as a cutoff. The transition from the XY phase to the VBS phase is further validated through level crossing (see Appendix B for details) the order parameter [c.f. Fig.3(b)] characterizing the VBS phase under open boundary conditions, as expressed in:
@@
As an example, we compute the fidelity susceptibility and the order parameters at $t_{2}/t_{1}=0.7$ [cf. Fig.5]. It is observed that the system resides in the VBS phase for $2.120<V/t_{1}<$ 2.12, characterized by a non-zero $O_{\text{VBS}}$ and vanishing $O_{\text{AFM}}$. Conversely, in the regime $32.12<V/t_{1}<$ 3, the reverse occurs, indicating that the system is in the AFM phase [cf. Fig.5(b)]. Notably, both order parameters, $O_{\text{AFM}}$ and $O_{\text{VBS}}$, seem to smoothly vanish at this same point, suggesting an unconventional nature for this quantum phase transition. This transition is argued to be a second-order phase transition with the correlation length critical exponent $\nu=1.11$ (see Appendix C for details), as demonstrated in Fig.5(a), (c), and (d) from the finite-size scaling of the fidelity susceptibility [52, 53, 54],
$\chi_{L}^{m}\propto L^{2/\nu-1}.$ (9)
where, $\chi_{L}^{m}$ represents the fidelity susceptibility per site at the peak position $\lambda=\lambda_{m}$, which converges towards the critical point $\lambda_{c}$ as $L$ tends to infinity. As the AFM phase breaks the $\mathbb{Z}_{2}$ spin-flip symmetry and the VBS phase breaks translation symmetry, the direct continuous second-order phase transition occurring between the AFM phase and the VBS state indicates that the phase transition is a DQCP beyond the LGW theory.
Interestingly, the DQCP arises from the merging of two BKT transitions. These transitions can be understood through the concept of domain walls, which are topological defects in one-dimensional systems. Beginning from the VBS phase (i.e at $t_{2}/t_{1}=0.7$ and $V=0$), characterized by, increasing the interaction strength $V/t_{1}$ induces the formation of AFM domain walls. Domain walls in VBS bind the AFM phase (and vice versa), leading to the destruction of the VBS phase and the emergence of the AFM phase as $t_{2}/t_{1}$ increases further. Consequently, a direct transition between the VBS phase to the AFM phase occurs under conditions of large $t_{2}$ and $V$, indicating the presence of a DQCP. This DQCP is linked to the two BKT transitions through a multicritical point characterized by an emergent high symmetry.
We note that our finding is analogous to the phenomena described in frustrated two-dimentional models [59, 60], where a gapless quantum spin liquid (QSL) phase gradually develops from the AFM-VBS transition, leading to two phase transitions between gapped and gapless phases (AFM-QSL transition and QSL-VBS transitions). Consequently, a DQCP emerging from two BKT transitions may be a universal phenomenon, offering crucial insights into the understanding of both BKT transitions and DQCPs.
@@
Appendix A Non-interacting energy spectrum
@@
$\begin{split}b_{r}=\frac{1}{\sqrt{N}}\sum e^{ikj}b_{k},\\ b_{r}^{\dagger}=\frac{1}{\sqrt{N}}\sum e^{-ik^{\prime}j}b_{k^{\prime}}^{\dagger},\end{split}$ (11)
the model described in Eq.(10) can be transformed into momentum space,
$\begin{split}H&=t_{1}\sum_{kk^{\prime}}\delta_{kk^{\prime}}e^{ik}b_{k^{\prime}}^{\dagger}b_{k}+t_{1}\sum_{kk^{\prime}}e^{-ik^{\prime}}b_{k^{\prime}}^{\dagger}b_{k}+t_{2}\sum_{kk^{\prime}}\delta_{kk^{\prime}}e^{2ik}b_{k^{\prime}}^{\dagger}b_{k}+t_{2}\sum_{kk^{\prime}}\delta_{kk^{\prime}}e^{-2ik^{\prime}}b_{k^{\prime}}^{\dagger}b_{k}\\ &=\sum_{k}\left[t_{1}(e^{ik}+e^{-ik})+t_{2}(e^{2ik}+e^{-2ik})\right]b_{k}^{\dagger}b_{k}\\ &=\sum_{k}2t_{1}\cos{k}+2t_{2}\cos{2k}b_{k}^{\dagger}b_{k}\\ &=\sum_{k}\epsilon(k)b_{k}^{\dagger}b_{k}\\ \end{split}$ (12)
Figure 6: The energy spectrum of free bosons with different $t_{2}/t_{1}$.
@@
$\epsilon(k)=2t_{1}(\cos{k}+j\cos{2k}),$ (13)
Here, $j={t_{2}}/{t_{1}}$. Using trigonometric relationships, $\cos{2k}=2\cos^{2}k+1$, the spectrum can be transformed as:
$\epsilon(k)=2t_{1}(2j\cos^{2}k+\cos{k}+j).$ (14)
@@
$\frac{d}{dk}\epsilon(k)=-2t_{1}(4j\cos{k}\sin{k}+\sin{k})=0$ (15)
The equation $\sin{k}=0$ or $01+4j\cos{k}=$ 0 offers two solutions, thereby affecting the spectrum $\epsilon(k)$ depending on the frustration parameter $j$ [24]. Specifically: When $j<\frac{1}{4}$, the dispersion $\epsilon(k)$ displays a single minimum at $k=\pi$ because $\sin{k}=0$. Conversely, when $j>\frac{1}{4}$, the spectrum $\epsilon(k)$ exhibits two distinct minima located at $k=\pm\arccos\left(-\frac{1}{4j}\right)$, arising from the condition $01+4j\cos{k}=$ 0. The critical value $j=\frac{1}{4}$ marks a special point known as the Lifshitz point. The system exhibits either a superfluid phase when $j<\frac{1}{4}$ or a chiral superfluid phase when $j>\frac{1}{4}$.
@@
Figure 7: The energy difference of the FDBH model with respect to the next-nearest-neighbor hopping. The red squares represent the energy difference between the ground state and the first excited state with $S_{total}^{z}=0$. The blue circles represent the energy difference between the ground state of $S_{total}^{z}=1$ and $S_{total}^{z}=0$. For $L=12$, the crossing point occurs at $t_{2c}$=0.37. For $L=20$, the crossing point locates at $t_{2c}=0.39$.
Appendix B Level crossing for the BKT transition
@@
$S_{\text{total}}^{z}=\sum_{r}{S_{r}^{z}}.$ (17)
When the number of particles in the system is a multiple of four, the ground state is a non-degenerate singlet state [49]. In the XY phase, the first excited state exhibits $S_{total}^{z}=1$, whereas in the VBS phase, the first excited state exhibits $S_{total}^{z}=0$. Therefore, the XY-VBS phase transition can be identified by the level crossing between the doublet and the singlet of the first excited states.
To identify the phase transition via level crossings, we investigated the energy gap $\Delta E=E_{m}^{(n)}-E_{m^{\prime}}^{(n^{\prime})}$ of the FDBH model using exact diagonalization with lattice sizes $L=12$ and $L=20$ at $V=0.1$ for $0.50<t_{2}/t_{1}<$ 0.5 as illustrated in Fig.7. Here, $E_{m}^{(n)}$ denotes energy levels where $m=0,\pm 1$ represents the total spin along the $z$-direction, and $n$ denotes the energy level (e.g., ground state, first excited state). The red squares depict the energy difference between the singlet ground state and the first excited state with $S_{total}^{z}=0$, which is non-degenerate except at the Majumdar-Ghosh point $t_{2}=0.5$. The blue circles correspond to the energy difference between ground states with $S_{total}^{z}=1$ and $S_{total}^{z}=0$. The critical point is identified where these two lines cross. For $L=12$, the XY phase to VBS phase transition occurs at $t_{2c}=0.37$, while for $L=20$, it shifts slightly to $t_{2c}=0.39$. This shift to higher $t_{2c}$ values with increasing lattice size indicates a pronounced finite-size effect.. In the hard-core boson limit, spin conservation is related to particle number conservation through the equation,
$S_{total}^{z}=L/2-N_{total},$ (18)
where $L$ is the total number of lattice sites and $N_{total}$ represents the total number of particles in the system. This mapping allows us to interpret spin states in terms of particle configurations. For instance, when $S_{total}^{z}=0$, the system corresponds to a half-filled particle case. By varying the particle number within a specific subspace, one can identify excited states.
Appendix C Correlation length critical exponent
@@
$\chi_{F}(\lambda_{m})\sim L^{2/\nu},$ (19)
for second-order phase transitions. Here, $\nu$ denotes the correlation length critical exponent, and $\lambda_{m}$ represents the peak position of the fidelity susceptibility for a system of size $L$. The correlation length critical exponent is determined by fitting the logarithm of both sides of Eq. (19):
$\ln\chi_{F}(\lambda_{m})\propto 2/\nu\ln L$ (20)
@@
Figure 8: The critical exponents of the correlation length for VBS-AFM transitions.
In the vicinity of the phase transition, we keep the interaction strength fixed and compute the ground state fidelity of the system across various sizes (including $L=128,144,192,240,\text{and}288$) as a function of the next-nearest-neighbor hopping $t_{2}/t_{1}$. By fitting the maximum fidelity against system size using the logarithmic relationship in Eq. (20), we extract the slope of the fit, thereby determining the correlation length critical exponent, as illustrated in Fig. 8. From the figure, it is observed that the critical exponent decreases gradually with increasing interaction strength. This suggests that the system exhibits varying universality classes for different quantum critical points.

### v2

@@ Abstract
In this letter, we instead explore a frustrated dipolar Bose-Hubbard (FDBH) model in one dimension under the hard-core limit, where nearest-neighbor interactions stem from dipolar interactions [36], which are achievable within an atomic mixture [35] as well. We show that the FDBH model hosts DQCPs, aligning with recent theoretical proposals [13, 37, 38, 39, 40] that describe a continuous quantum phase transition from an antiferromagnetic (AFM) state to a valence bond solid (VBS) state. We employ the density matrix renormalization group (DMRG) technique [41, 42, 43, 44, 45] to investigate quantum phase transitions of FDBH model by varying nearest-neighbor interactions. Our study reveals the presence of three distinct phases, delineated by two BKT transitions and one DQCP through the fidelity susceptibility. Intriguingly, our proposal reveals that DQCPs emerge from the fusion of two BKT transitions, providing an essential insight into the understanding of both BKT transitions and DQCPs.
@@
by utilizing the transformations $\sigma^{x}_{r}=b^{\dagger}_{r}+b_{r},~{}~{}\sigma^{y}_{r}=i(b^{\dagger}_{r}-b_{r}),~{}~{}\sigma^{z}_{r}=1-2n_{r}$, with $(\sigma^{x}_{r},\sigma^{y}_{r},\sigma^{z}_{r})$ being Pauli matrices. Here, the particle occupation numbers are represented by spins.
@@
When $V=0$, the system shown in Eq.(3) become the XY model incorporating both the nearest-neighbor and next-nearest-neighbor hopping terms. In particular, when $t_{2}=0$, the system’s ground state is characterized by the gapless XY (superfluidity (SF) in bosonic language) phase (see Supplemental Material [46] for details). Upon increasing $t_{2}$, the system undergoes a phase transition from the XY phase to the VBS (bond-ordered-wave (BOW) in bosonic language) phase [35, 47, 48, 49, 50]. To verify if the transition between the XY phase and the VBS phase persists in the presence of interaction $V$, we calculate the ground-state phase diagram of the FDBH model described in Eq.(2) varying parameters $t_{2}/t_{1}$ and $V/t_{1}$ in terms of DMRG method. During DMRG simulations, we keep 500 matrix states and enforce $n_{\text{max}}=1$ per site. The full phase diagram is shown in Fig.2, which is obtained from the fidelity susceptibility [51, 52, 53, 54, 55, 56, 57, 58],
@@
where $\chi_{\infty}$ represents the fidelity susceptibility in the thermodynamic limit, $\lambda_{c}$ is the critical point, and $a$ is a positive nonuniversal constant serving as a cutoff. The transition from the XY phase to the VBS phase is further validated through level crossing (see Supplemental Material [46] for details) and the order parameter [c.f. Fig.3(b)] characterizing the VBS phase under open boundary conditions, as expressed in:
@@
As an example, we compute the fidelity susceptibility and the order parameters at $t_{2}/t_{1}=0.7$ [cf. Fig.5]. It is observed that the system resides in the VBS phase for $2.120<V/t_{1}<$ 2.12, characterized by a non-zero $O_{\text{VBS}}$ and vanishing $O_{\text{AFM}}$. Conversely, in the regime $32.12<V/t_{1}<$ 3, the reverse occurs, indicating that the system is in the AFM phase [cf. Fig.5(b)]. Notably, both order parameters, $O_{\text{AFM}}$ and $O_{\text{VBS}}$, seem to smoothly vanish at this same point, suggesting an unconventional nature for this quantum phase transition. This transition is argued to be a second-order phase transition with the correlation length critical exponent $\nu=1.11$ (see Supplemental Material [46] for additional critical exponents), as demonstrated in Fig.5(a), (c), and (d) from the finite-size scaling of the fidelity susceptibility [53, 54, 55],
$\chi_{L}^{m}\propto L^{2/\nu-1},$ (9)
where $\chi_{L}^{m}$ represents the fidelity susceptibility per site at the peak position $\lambda=\lambda_{m}$, which converges towards the critical point $\lambda_{c}$ as $L$ tends to infinity. As the AFM phase breaks the $\mathbb{Z}_{2}$ spin-flip symmetry and the VBS phase breaks translation symmetry, the direct continuous second-order phase transition occurring between the AFM phase and the VBS state indicates that the phase transition is a DQCP beyond the LGW theory.
Interestingly, the DQCP arises from the merging of two BKT transitions. These transitions can be understood through the concept of domain walls, which are topological defects in one-dimensional systems. Beginning from the VBS phase (i.e at $t_{2}/t_{1}=0.7$ and $V=0$), increasing the interaction strength $V/t_{1}$ induces the formation of AFM domain walls. Domain walls in VBS bind the AFM phase (and vice versa), leading to the destruction of the VBS phase and the emergence of the AFM phase as $V/t_{1}$ increases further. Consequently, a direct transition between the VBS phase to the AFM phase occurs under conditions of large $t_{2}/t_{1}$ and $V/t_{1}$, indicating the presence of a DQCP. This DQCP is linked to the two BKT transitions through a multicritical point characterized by an emergent high symmetry.
We note that our finding is analogous to the phenomena described in frustrated two dimentional models [60, 61], where a gapless quantum spin liquid (QSL) phase gradually develops from the AFM-VBS transition, leading to two phase transitions between gapped and gapless phases (AFM-QSL transition and QSL-VBS transitions). Consequently, a DQCP emerging from two BKT transitions may be a universal phenomenon, offering crucial insights into the understanding of both BKT transitions and DQCPs.
@@
Supplemental Material
We present additional information on (1) Non-interacting energy spectrum, (2) Level crossing for the BKT transition and (3) Correlation length critical exponent.
A: Non-interacting energy spectrum
@@
$\begin{split}b_{r}=\frac{1}{\sqrt{N}}\sum e^{ikj}b_{k},\\ b_{r}^{\dagger}=\frac{1}{\sqrt{N}}\sum e^{-ik^{\prime}j}b_{k^{\prime}}^{\dagger},\end{split}$ (S2)
the model described in Eq.(S1) can be transformed into momentum space,
$\begin{split}H&=t_{1}\sum_{kk^{\prime}}\delta_{kk^{\prime}}e^{ik}b_{k^{\prime}}^{\dagger}b_{k}+t_{1}\sum_{kk^{\prime}}\delta_{kk^{\prime}}e^{-ik^{\prime}}b_{k^{\prime}}^{\dagger}b_{k}+t_{2}\sum_{kk^{\prime}}\delta_{kk^{\prime}}e^{2ik}b_{k^{\prime}}^{\dagger}b_{k}+t_{2}\sum_{kk^{\prime}}\delta_{kk^{\prime}}e^{-2ik^{\prime}}b_{k^{\prime}}^{\dagger}b_{k}\\ &=\sum_{k}\left[t_{1}(e^{ik}+e^{-ik})+t_{2}(e^{2ik}+e^{-2ik})\right]b_{k}^{\dagger}b_{k}\\ &=\sum_{k}\left(2t_{1}\cos{k}+2t_{2}\cos{2k}\right)b_{k}^{\dagger}b_{k}\\ &=\sum_{k}\epsilon(k)b_{k}^{\dagger}b_{k}\\ \end{split}$ (S3)
Figure S1: The energy spectrum of free bosons with different $t_{2}/t_{1}$.
@@
$\epsilon(k)=2t_{1}(\cos{k}+j\cos{2k}),$ (S4)
Here, $j={t_{2}}/{t_{1}}$. Using trigonometric relationships, $\cos{2k}=2\cos^{2}k-1$, the spectrum can be transformed as:
$\epsilon(k)=2t_{1}(2j\cos^{2}k+\cos{k}-j).$ (S5)
@@
$\frac{d}{dk}\epsilon(k)=-2t_{1}(4j\cos{k}\sin{k}+\sin{k})=0.$ (S6)
The equations $\sin{k}=0$ and $01+4j\cos{k}=$ 0 offer two solutions, thereby affecting the spectrum $\epsilon(k)$ depending on the frustration parameter $j$ [24]. Specifically: When $j<\frac{1}{4}$, the dispersion $\epsilon(k)$ displays a single minimum at $k=\pi$ because $\sin{k}=0$. Conversely, when $j>\frac{1}{4}$, the spectrum $\epsilon(k)$ exhibits two distinct minima located at $k=\pm\arccos\left(-\frac{1}{4j}\right)$, arising from the condition $01+4j\cos{k}=$ 0. The critical value $j=\frac{1}{4}$ marks a special point known as the Lifshitz point. The system exhibits either a superfluid phase when $j<\frac{1}{4}$ or a chiral superfluid phase when $j>\frac{1}{4}$.
@@
Figure S2: The energy difference of the FDBH model with respect to the next-nearest-neighbor hopping. The red squares represent the energy difference between the ground state and the first excited state with $S_{\text{total}}^{z}=0$. The blue circles represent the energy difference between the ground state of $S_{\text{total}}^{z}=1$ and $S_{\text{total}}^{z}=0$. For $L=12$, the crossing point occurs at $t_{2c}$=0.37. For $L=20$, the crossing point locates at $t_{2c}=0.39$.
B: Level crossing for the BKT transition
@@
$S_{\text{total}}^{z}=\sum_{r}{S_{r}^{z}}.$ (S8)
When the number of particles in the system is a multiple of four, the ground state is a non-degenerate singlet state [50]. In the XY phase, the first excited state exhibits $S_{\text{total}}^{z}=1$, whereas in the VBS phase, the first excited state exhibits $S_{\text{\text{total}}}^{z}=0$. Therefore, the XY-VBS phase transition can be identified by the level crossing between the doublet and the singlet of the first excited states.
To identify the phase transition via level crossings, we investigated the energy gap $\Delta E=E_{m}^{(n)}-E_{m^{\prime}}^{(n^{\prime})}$ of the FDBH model using exact diagonalization with lattice sizes $L=12$ and $L=20$ at $V=0.1$ for $0.50<t_{2}/t_{1}<$ 0.5 as illustrated in Fig.S2. Here, $E_{m}^{(n)}$ denotes energy levels where $m=0,\pm 1$ represents the total spin along the $z$-direction, and $n$ denotes the energy level (e.g., ground state, first excited state). The red squares depict the energy difference between the singlet ground state and the first excited state with $S_{\text{total}}^{z}=0$, which is non-degenerate except at the Majumdar-Ghosh point $t_{2}=0.5$. The blue circles correspond to the energy difference between ground states with $S_{\text{total}}^{z}=1$ and $S_{\text{total}}^{z}=0$. The critical point is identified where these two lines cross. For $L=12$, the XY phase to VBS phase transition occurs at $t_{2c}=0.37$, while for $L=20$, it shifts slightly to $t_{2c}=0.39$. This shift to higher $t_{2c}$ values with increasing lattice size indicates a pronounced finite-size effect. In the hard-core boson limit, spin conservation is related to particle number conservation through the equation,
$S_{\text{total}}^{z}=L/2-N_{\text{total}},$ (S9)
where $L$ is the total number of lattice sites and $N_{\text{total}}$ represents the total number of particles in the system. This mapping allows us to interpret spin states in terms of particle configurations. For instance, when $S_{\text{total}}^{z}=0$, the system corresponds to a half-filled particle case. By varying the particle number within a specific subspace, one can identify excited states.
Figure S3: The critical exponents of the correlation length for VBS-AFM transitions.
C: Correlation length critical exponent
@@
$\chi_{F}(\lambda_{m})\sim L^{2/\nu},$ (S10)
for second-order phase transitions. Here, $\nu$ denotes the correlation length critical exponent, and $\lambda_{m}$ represents the peak position of the fidelity susceptibility for a system of size $L$. The correlation length critical exponent is determined by fitting the logarithm of both sides of Eq. (S10):
$\ln\chi_{F}(\lambda_{m})\propto 2/\nu\ln L.$ (S11)
@@
In the vicinity of the phase transition, we keep the interaction strength fixed and compute the ground state fidelity susceptibility of the system across various sizes (including $L=128,144,192,240$, and $288$) as a function of the next-nearest-neighbor hopping $t_{2}/t_{1}$. By fitting the maximum fidelity against system size using the logarithmic relationship in Eq. (S11), we extract the slope of the fit, thereby determining the correlation length critical exponent, as illustrated in Fig. S3. From the figure, it is observed that the critical exponents decrease gradually with increasing interaction strength. This suggests that the system exhibits varying universality classes for different quantum critical points.